Divide this fraction
6p - 18
over 9p
divided by
3r - 9
over p^2 + 2p
Should I be simplfying first or is this already done. I need help on this one.
Thanks

Answer by fractalier(2101)

From
[(6p - 18)/9p] / [(3p - 9)/(p^2 + 2p)] =
we factor and cancel giving us
[6(p - 3) / 9p] / [3(p - 3) / p(p+2)] =
now invert and multiply
[6(p - 3) / 9p] * [p(p+2) / 3(p - 3)] =
now cancel and reduce where we can...
(2 / 3p) * [p(p+2) / 3] =
2(p+2) / 9